Reports 1-1 of 1 Clear search Modify search
AdV-ISC (Commissioning up to first full interferometer lock)
bonnand - 13:28 Thursday 02 December 2021 (54075) Print this report
Comment to ISC shift: SDB1 drift control, long lock (54067)

Something worth noticing concerning the long lock of yesterday is that the cavity power was much lower than the previous attempts.

Figure shows the cavity power (B7/B8_DC) and the B4_12_MHz and B4_112MHz.

One can see that at the begining of the lock the power in the cavity follows the usual trend but after a few minutes dropped at 13:13 UTC and become noisier and then keep lowering up to the moment where the NE_Ty was moved (not visible in the plot).

I don't know what causes the cavity power to drop but it is not related to the opening of the galvo loop that happened 9 minutes later at around 13:22.

It is also noticeable that the B4 12MHz and 112MHz magnitude is increasing while the cavity power is decreasing but I cannot say if it is related.


Images attached to this comment
Search Help